You have decided to refinance your mortgage. You plan to borrow whatever is outstanding on your current mortgage. The current monthly payment is $1,912 and you have made every payment on time. The original term of the mortgage was 30 years, and the mortgage is exactly for years and eight months old. You have just made your monthly payment. The mortgage interest rate is 5.507% (APR). How much do you owe on the mortgage today?
Answer #1

B = (PMT/R) x (1 - (1/(1+R)^N) where B = Balance Principle, PMT= monthly installment, R= APR and N= remaining installments

B = (1912/(0.05507/12)) x (1 - (1/(1+(0.05507/12))^(360-8) ) = 1912/0.004589166 * ( 1- 1/(1.004589166)^352 )

B = 1912/0.004589166 * ( 1- 1/5.011278615 ) = $333494.30

We owe on the mortgage today = $333494.30